OpenStudy (anonymous):

Simplify 3 square root of 5 end root minus 2 square root of 7 end root plus square root of 45 end root minus square root of 28 . 2 square root of 12 2 square root of 2 6 square root of 5 end root minus 4 square root of 7 6 square root of 10 end root minus 4 square root of 14

OpenStudy (anonymous):

@Directrix can you help?

OpenStudy (calculusxy):

\[3\sqrt{5} - 2\sqrt{7} + \sqrt{45}- \sqrt{28}\] Finding the answer to these problems is almost like combining like terms in regular expressions. However, in this case you must have the same number under the radical sign. For example: \[2\sqrt{7} + 3\sqrt{7} = 5\sqrt{7}\] But you must first try to figure to fully simplify \(\sqrt{45}\) and \(\sqrt{28}\) and then start to combine.

OpenStudy (anonymous):

Im confused...

OpenStudy (anonymous):

@calculusxy

OpenStudy (anonymous):

you can only add and subtract if the number under the radical is the same, so you need to reduce \(\sqrt{45}\) and \(\sqrt{28}\) first. for example\[\sqrt{45}=\sqrt{9\times5}=\sqrt{9}\times \sqrt{5}=3\sqrt{5}\] try with \(\sqrt{28}\)